OpenStudy (anonymous):

10x+9y=21 2x-18y=24 Solve Each System By Elimination And Steps Only Thanks .

OpenStudy (anonymous):

Solve: 10x+9y=21 Step 1: Add -9y to both sides. 10x+9y+-9y=21+-9y 10x+-9y+21 Divide by 10 on both sides

OpenStudy (anonymous):

10x/10=-9y+21/10 x=-9y/10+21/10

OpenStudy (anonymous):

Solve:2x-18y=24 add 18y to both sides: 2x-18y+18y=24+18y 2x+18y+24 Divide both sides by two: 2x/2=18y+24/2 x=9y+12

OpenStudy (anonymous):

reply if I helped

OpenStudy (anonymous):

You Helped !!!!!!!!!!!! Thanks !

OpenStudy (anonymous):

u are so welcome :)